OpenStudy (anonymous):

Solve the equation 2 over 3 (x − 2) = 4x for x. negative 2 over 5 negative 5 over 2 2 over 5 5

OpenStudy (campbell_st):

if the problem is \[\frac{2}{3} (x - 2) = 4x\] start by multiplying both sides of the equation by 3, which eliminates the fraction \[2(x -2) = 12x\] divide both sides of the equation by 2 x - 2 = 6x now you can solve for x. hope this helps